solve for x and y from given equations
3 to the power 3x+1=9 to the power y and x squared+2y=29

Answer by t0hierry(194)

3 power 3x+1 = 9 power y = 3 power 2y
3x + 1 = 2y
x^2 + 2y = 29
substituting gives
x^2 + 3x + 1 = 29
x=4